OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. x^5/3 = 243

hero (hero):

x^5 = 243*3

hero (hero):

5th root(x^5) = 5th root{729}

hero (hero):

x = 3.73719

hero (hero):

I don't see a way to check for extraneous solutions here